OpenStudy (anonymous):

write an equation of the line containing the given point and paraelle to the given line express answer in the y=mx+b form (-6,6);7x=6y+5

OpenStudy (anonymous):

Put (-6, 6) in\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

\[7x=6y+5\] \[6y=7x-5\]\ \[y=\frac{7}{6}x-\frac{5}{6}\] gives the slope of the line as \[\frac{7}{6}\] so your line must have slope \[\frac{7}{6}\] also. the point is \[(-6,6)\] so use the almighty point slope formula to find the equation for your line. \[y-y_1=m(x-x_1)\] \[y-6=\frac{7}{6}(x-(-6))\]

OpenStudy (anonymous):

\[y-6=\frac{7}{6}(x+6)\] \[y-6=\frac{7}{6}x+7\] \[y=\frac{7}{6}x+13\]